Circular No. 6805

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 980109 J. in ’t Zand and J. M. Muller, Space Research Organisation, Utrecht;

L. Piro, L. Di Ciolo, A. Coletta, S. Rebecchi, and G. Gennaro, BeppoSAX Science Operation Center, Telespazio, Rome; M. Feroci, Istituto di Astrofisica Spaziale, CNR, Rome; and F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, report on behalf of the BeppoSAX team: “A γ-ray burst (Bacodine trigger 6564) has been detected by the BeppoSAX Gamma-Ray burst Monitor on Jan. 9.05024 UT. A quicklook analysis shows a single-peak γ-ray burst lasting ∼ 25 s with a peak intensity of 550 counts/s. The γ-ray burst was also detected by the Wide Field Camera WFC2 with a peak flux of ∼ 0.8 Crab. The position given by an off-line analysis is α = 0^{h}25^{m}56^{s}, δ = −63^{o}01′.4 (equinox 2000.0). The event was detected in a part of the orbit when the satellite was in a very unfavorable attitude control configuration, resulting in a conservative 99-percent error radius of 10′. A BeppoSAX Narrow Field follow-up observation will not be performed. Observations at all wavelengths are solicited.”

SUPERNOVA 1998A IN IC 2627 Andrew Williams, Perth Observatory, reports the discovery by Simon

Woodings, Ralph Martin, and himself of an apparent supernova (R ∼ 17) on a CCD frame taken on Jan. 6.77 UT with the 0.61-m Perth-Lowell reflector as part of the Perth Astronomy Research Group’s automated supernova search. SN 1998A is located 40′′ east and 10′′ south of the center of IC 2627. No object was present at this position on a frame taken on 1997 Dec. 18 (limiting mag R ∼ 18.5), nor is any star present on earlier, deeper search images of this galaxy. A. Verveer and J. Biggs obtained an unfiltered CCD image on Jan. 7.844 with the Perth 0.25-m Mike Candy Telescope, from which Biggs has derived the following precise position for SN 1998A: α = 11^{h}09^{m}50^{s}.33 ± 0^{s}.07, δ = −23^{o}43′43′′.1 ± 0′′.2 (equinox 2000.0; GSC reference stars); the magnitude was estimated as 16.5.

SUPERNOVA 1997ei IN NGC 3963 S. Moretti and S. Tomaselli, Forlí, Italy, report the following unfiltered

CCD magnitude for SN 1997ei: Jan. 8.85 UT, 15.9 (GSC reference stars).

1998 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green